nun con Un in normado 3. Assuming that X ~ U[-1,1], derive the density of Y := - In(X + 1). 211 1 . L . TIC
Answer #1

It is given that Xu u (-2, LJ then pdf of x will be fra L 15XS? the IT-7-1). otherwise -16 n? I げた otherwise cs Scanned with1 here y is defined as YzIen(xb) let us find out the calf of y le & Flil Plycy IP [-en(x+1) 5 y] P[ln (X+1)-4) - P[-x+12_ey =so pdf -flyl of y d f(y= ley sinu-L < X LI andy = -In (x+2) L - y e > ln{X+1) Ye XF [ x = e) L LSX EL - ا ر لا ocele 2 = 2 >